OpenStudy (anonymous):

Using the given points, determine the slope. (8, 1) and (8, -4)

OpenStudy (anonymous):

Easy way to find the slope is: Tell me first what is the y difference between the two points?? Means to say that Difference of y coordinates is??? Tell me..

OpenStudy (anonymous):

5

OpenStudy (anonymous):

And now the difference of x coordinates is ???

OpenStudy (anonymous):

0

OpenStudy (anonymous):

m=(y1-y2)/(x1-x2)=(1--4)/(8-8)=5/0 , thus impossible or inf

OpenStudy (anonymous):

Just divide them: Difference of y coordinates/Difference of x coordinates.. What you get??